OpenStudy (anonymous):

Is it possible to graph the equation y = -(6/7)x + 6?

OpenStudy (a_clan):

yes 6x+7y=42

OpenStudy (anonymous):

without it going below the x axis, I should add..

OpenStudy (amistre64):

i dunno, that might be one of those not doable lines ....

OpenStudy (amistre64):

anything with a slope other than zero will go below the x axis at some point

OpenStudy (anonymous):

It involves a story problem though...so I don't think 'no solution' is a choice...

OpenStudy (anonymous):

Did you re-write it in the y= mx + b form? It has a portion of the line in the +/+ region.

OpenStudy (anonymous):

I rewrote 6x + 7y = 42 as y = -(6/7)x + 6..

OpenStudy (anonymous):

Indeed. So the y intercept is?

OpenStudy (anonymous):

y intercept is 6...but the negative 6 over 7 is throwing me off, as far as graphing the line..

OpenStudy (a_clan):

Any line can be extended in either direction. So, line has to go below X axis

OpenStudy (anonymous):

It's ok. so that means you will go down 6 and over 7.

OpenStudy (anonymous):

So there's still quite a bit of area above the x axis.

OpenStudy (anonymous):

So I just draw the one point directly on the x axis...and draw the line between the two points?

OpenStudy (anonymous):

Yep.

OpenStudy (a_clan):

If you have problem graphing slope-intercept form, convert it into standard form Ax+By=C. It is easier to graph

OpenStudy (anonymous):

the one involving the negative 6 over 7, that is..

OpenStudy (anonymous):

ok, then...

OpenStudy (anonymous):

You have the point (0,6) and (7,0)

OpenStudy (anonymous):

A Mathematica plot is attached.

OpenStudy (anonymous):

and all the points on the line between them.

OpenStudy (anonymous):

ok then...thanks a lot guys...it helps me to discuss out my math problems..

OpenStudy (anonymous):

of course!